we started with 5 ounces of perfume that was 95% alcohol. we added some preservative that was only 75% alcohol. we ended up with a perfume that was 91% alcohol. determine how many ounces of the preservative were added.
Answer by srinivas.g(540)

5 ounces of perfume consist of 95% alcohol .
so , intital quantity of alcohol = 0.95 * 5
=4.75 onces
lets assume x ounces of some preservative has been added
now alcohol in the preservative = 75% of x
=0.75 x
total quantity of alcohol = 4.75 +0.75 x
but total volume of perfume after adding preservative = 5+x
alcohol in the total quantity = 91 % of (5+x)
= 0.91*5+0.91 x
=4.55 +0.91 x

so 4.75 +0.75 x =4.55 +0.91 x
4.75-4.55 =0.91 x -0.75 x
0.2 = 0.16 x
divide with 0.16 on both sides
0.2/0.16 = 0.16 x/0.16
1.25 = x
x= 1.25 onces
so amount of preservative added = 1.25 ounces